# Leveraging Sparse Annotations for Leukemia Diagnosis on the Large Leukemia Dataset

### v1

@@ Abstract
Leukemia is $10^{th}$ most frequently diagnosed cancer and one of the leading causes of cancer-related deaths worldwide. Realistic analysis of Leukemia requires White Blook Cells (WBC) localization, classification, and morphological assessment. Despite deep learning advances in medical imaging, leukemia analysis lacks a large, diverse multi-task dataset, while existing small datasets lack domain diversity, limiting real-world applicability. To overcome dataset challenges, we present a large-scale WBC dataset named ‘Large Leukemia Dataset’ (LLD) and novel methods for detecting WBC with their attributes. Our contribution here is threefold. First, we present a large-scale Leukemia dataset collected through Peripheral Blood Films (PBF) from several patients, through multiple microscopes, multi-cameras, and multi-magnification. To enhance diagnosis explainability and medical expert acceptance, each leukemia cell is annotated at 100x with 7 morphological attributes, ranging from Cell Size to Nuclear Shape. Secondly, we propose a multi-task model that not only detects WBCs but also predicts their attributes, providing an interpretable and clinically meaningful solution. Third, we propose a method for WBC detection with attribute analysis using sparse annotations. This approach reduces the annotation burden on hematologists, requiring them to mark only a small area within the field of view. Our method enables the model to leverage the entire field of view rather than just the annotated regions, enhancing learning efficiency and diagnostic accuracy. From diagnosis explainability to overcoming domain-shift challenges, presented datasets could be used for many challenging aspects of microscopic image analysis. The datasets, code, and demo are available at: https://im.itu.edu.pk/sparse-leukemiaattri/.
@@
Explainable diagnosis, Blood cancer, leukemia, attribute, spare annotation, multi-task learning
@@ 1 Introduction
Leukemia is among the leading causes of cancer-related deaths worldwide, with an estimated 487,000 new cases and 305,000 deaths reported in 2022 [1]. This disease disproportionately affects people under the age of 40, with children being particularly impacted [2]. The diagnosis of Leukemia requires an expert hematologist with extensive experience, capable of analyzing peripheral blood slides employing costly medical equipment. Given the prerequisites for the diagnosis and prognosis of leukemia, many remote regions in underdeveloped countries experienced higher mortality rates [3]. However, early detection of leukemia enables timely and accurate treatment, which can significantly improve survival rates. For leukemia diagnosis, hematologists examine the Peripheral Blood Film (PBF) of patients a process that is often tedious, error-prone, and requires significant expertise. The limited availability of experts and costly diagnostic resources hinders early, precise, and timely detection of leukemia, particularly in resource-constrained areas. Furthermore, leukemia treatment requires repetitive blood sample preparation and detailed analysis requiring frequent access to experts.
AI-driven solutions are vital to overcome the low accessibility of experts. Such a system should be able to work across multiple microscopes, and digitization scenarios, and to augment hematologists, it should provide explainable information to hematologists. Unfortunately, existing leukemia datasets  [4, 5, 6, 7, 8, 9] suffer from many limitations, including the absence of localization annotations of white blood cells, the lack of associated morphological attributes, and a limited number of images [10]. Furthermore, there is no established multi-task learning (MTL) [11, 12, 13] approach that simultaneously detects white blood cells and predicts their morphological attributes.
@@
In addition to these limitations, these supervised algorithms ([14, 15, 16, 17, 18]) require every cell in each image to be annotated, adding a significant manual burden. In the practical scenario, a hematologist uses only a small portion of the acquired image to perform the required analysis and annotation. Thus majority of the cells remain unannotated. On the other hand, requiring annotation of complete blood films not only increases time and cost but also limits the number of blood films that could be annotated [^1] Note that multiple image patches are extracted at the specified resolution. Employing existing object detection strategies over sparsely annotated datasets results in subpar performance since these methods are unable to use unlabeled parts of slides.Complete annotation of a blood film in the monolayer takes around 200 minutes—eight times longer than the 25 minutes needed for sparse annotation.
@@
To overcome dataset challenges, we present a large-scale WBC dataset named ‘Large Leukemia Dataset’ (LLD), that consists of two subsets ‘LeukemiaAttri’ [^2] This subset was published in ‘A Large-scale Multi Domain Leukemia Dataset for the WBC Detection with Morphological Attributes for Explainability’ at MICCAI 2024 [19], and this journal paper serves as its extension., capturing cell localization, classification, and morphological attribute prediction tasks; and ‘Sparse-LeukemiaAttri’, for the object-detection and attribute prediction based on sparsely annotated dataset task. An important highlight of the LeukemiaAttri  is that it has been collected through two microscopes from different price ranges, one high-cost (HCM) and one low-cost (LCM), at three magnifications (100x, 40x, 10x) using various sensors: a high-end camera for HCM, a mid-range camera for LCM, and a mobile phone camera for both microscopes. In this collection, using high-resolution HCM (100x), experienced hematologists annotated 10.3k WBC of 14 types, having 55k morphological labels from 2.4k images of several PBS leukemia patients. The data set collection procedure is illustrated in Fig.2 and further details are in section 3.
@@
In conjunction with these datasets, we propose two models named Leukemia Attribute Detector (AttriDet) and Sparse Leukemia Attribute Detector (SLA-Det). For the task of detection of white blood cells (WBCs) along with the prediction of morphological attributes, we integrated a lightweight head, named AttriHead, into the existing object detector. AttriHead employs multiscale features from the feature backbone used in object detection. This not only enhances computational efficiency but also enables the backbone to become more informative through MTL. For the sparsely-supervised object detection, we propose a novel algorithm called SLA-Det. Note that in our case on average only 20% of the image is annotated shown in Figure 1. To effectively use the unannotated region information, SLA-Det incorporates region-guided masking and pseudo-label selection based on multiple criteria, ensuring robust model performance. We believe that these extensions align our dataset and framework more with clinical practices and enhance the reliability of automated leukemia detection. Our extensive experimental results and discussion validate the proposed ideas and framework. Our contributions are given below.
@@
    Key results: Using sparse annotations, SLA-Det can process slides at 40x resolutions while maintaining accuracy, unlike AttriDet, which only worked on 100x resolution; thus, it can process slides 4x faster than AtriDet. SLA-Det outperforms state-of-the-art sparse detection methods by 34.7 $mAP@_{50}$, demonstrating superior efficiency and performance.
@@ 2 Related Work
Supervised WBC detection: Deep learning (DL) techniques have been extensively applied to WBC classification [20, 21, 22, 23, 24]. Huang et al. [25] implemented a two-step approach for their end-to-end solution: first, detecting the WBCs, and then performing classification after cropping the detected cells. To classify single-cell images, Saidani et al. [26] optimized a CNN model and employed multiple pre-processing techniques, such as color conversion and image histogram equalization. Similarly, the authors in [27] classified two types of WBCs from the ALL-IDB data set [27]. G Liang et al. [28] developed a combined model of convolutional neural networks (CNN) and recurrent neural networks (RNN) to classify four types of WBCs using the BCCD dataset [29].
To detect WBCs in whole microscopy images, the authors in  [30] gave a lighter deep-learning CNN model and adapted the YOLOX-nano [31] baseline to identify the types of WBC. Their method enhances the overall model architecture by improving Intersection over Union (IoU) loss. A blood film and bone marrow aspirates-based multiple instance learning for leukocyte identification approaches have been proposed to classify different types of Leukemia  [32]. For image preprocessing, they mainly applied threshold methods.
@@
Sparsely annotated object detection: To tackle the challenge of object detection in sparsely annotated datasets, authors in [33] proposed a method that integrates supervised and self-supervised learning into their framework. The approach involves extracting features from an image and its multiple augmented versions, which are then processed through a shared region proposal network. A positive pseudo-mining technique is introduced to effectively filter out labeled, unlabeled, and foreground regions. This results in separate losses for labeled regions—supervised loss and self-supervised loss—ensuring consistent feature extraction across both original and augmented views. This combined method enhances the model’s generalization ability and improves object detection performance. Similarly, Z. Wu et al. [34] used a soft sampling technique to mitigate the impact of missing annotations. Their approach applies the Gompertz function to adjust gradient weights based on the overlap between positive instances and regions of interest. Niitani et al. [35] propose a part-aware learning approach that trains a model twice, using pseudo labels from the first model to guide the second. By leveraging spatial relationships, it selectively ignores classification loss for part categories within subject categories, improving training and performance with sparse annotations. To generate labels for sparsely annotated datasets, the co-mining approach is applied by authors of [36]. It involves constructing a Siamese detection network with two branches, each generating pseudo-label sets using a co-generation module. It is subsequently used as guided signals to the model, leveraging augmentation techniques to generate diverse pseudo-label sets. Finally, a co-generation module converts predictions from one branch into pseudo-labels for the other, facilitating complementary supervision between the network branches. Note that previous benchmarks, such as [33, 36], implemented sparse object detection algorithms on self-created datasets. In contrast, our dataset consists of small patches annotated with detailed morphological features and patch localization. In addition, we introduce a multi-headed WBC detection method tailored for such data.
@@ 3 Dataset
Leukemia comprises four main types: Acute lymphocytic leukemia (ALL), Chronic lymphocytic leukemia (CLL), Acute myeloid meukemia (AML), and Chronic myeloid meukemia (CML) [1]. These types are categorized into two major lineages: myeloid (AML and CML) and lymphoid (ALL and CLL). Each leukemia type exhibits distinct abnormalities, including increased myelo-blasts in CML, atypical lymphocytes in CLL, elevated myeloblasts in AML, and excessive lymphoblasts in ALL. However, as shown in Table 1, most existing white blood cell (WBC) datasets were collected using a single microscope, lacked bounding boxes and morphological details, and often included either only healthy individuals or patients with a single leukemia type.
@@
Multi. Multi. Cells BBX Multi. No. of WBC Morphology
Dataset Type Micro. in image Res. WBC’s Classes
IDB [4] ALL 510 (LB) 2
IDB2 [5] ALL 260 2
LISC [6] Normal 250 5
Munich[7] AML 18,365 15
Raabin [8] Normal 17,965 5
HRLS [9] Multi. 16,027 9
WBCAtt [37] Normal 10,298 5
@@ 3.1 Limitations of existing WBC datasets
The details of publicly available datasets on WBC and leukemia are shown in Table 1, which highlights various limitations such as: 1) The ALL-IDB dataset [4] contains only 510 WBCs from patients with Acute Lymphocytic Leukemia (ALL) and includes only cell centroid information. 2) In version 2 of the ALL IDB [5], each image contains only a single WBC, with the cell’s status labeled accordingly. 3) The LISC [6] dataset contains a smaller number of WBCs with five classes such as basophil, eosinophil, lymphocyte, monocyte, and neutrophil. 4) The entire dataset of Munich AML’s [7] contains single-cell images that come under the umbrella of AML. 5) Only Raabin [8], out of all the aforementioned existing datasets, has complete bounding box-level annotation with single-cell images, but the dataset consists entirely of normal individuals, with over 60% of the images containing only a single cell in the FoV. Furthermore, the Raabin-WBC collection contains images from two separate cameras and two different microscopes, without maintaining any correspondence across microscopes. 6) The HRLS [9] dataset is captured using a single microscope and covers only two types of Leukemia and nine types of WBCs. 7) WBCAtt dataset [37] is the improved version of [38], where annotations of attributes are added for each WBC. The dataset [38] lacks the abnormal WBCs and is captured using a single microscope only.
Figure 2: a) Illustrates the image acquisition procedure utilizing standard mobile camera (C1), premium camera for high-cost microscopy (HCM), and mid-range camera (C2) for low-cost microscopy (LCM) at various resolutions (100x, 40x, and 10x). The images are obtained with both high and low-cost microscopes. b) Displays our microscope set up with four cameras C1 to C4. C1 and C2 capture the field of view of the microscope lens, and C3 and C4 capture stage scale. c) This table presents various types of white blood cells (WBC) and their distinctive morphological features.
@@ 3.2 Our Dataset Collection Procedure
To understand leukemia prognosis, we consulted healthcare professionals to finalize key WBC types and their morphology. To improve deep learning across modalities, we propose capturing images of four leukemia PBF types from multiple patients at various magnifications using available resources. Capturing images at varying resolutions enhances deep learning generalization, enabling low-cost, automated leukemia diagnostics.
In this dataset collection, PBFs are sourced from a diagnostic lab with patient consent and include annotations from a hematologist in the monolayer area. Images are captured using two microscopes: the high-cost Olympus CX23 and the low-cost XSZ-107BN, along with three cameras: the HD1500T (HCM), ZZCAT 5MP (LCM), and the Honor 9x Lite mobile camera. The high-end camera (ZZCAT 5MP) is 47 times more expensive than the low-end camera (XSZ-107BN), while the HCM is 17 times more costly than the LCM. Although the LCM is more affordable, it sacrifices visual detail and has a smaller field of view. Note that capturing the same WBC across multiple microscopes and resolutions is highly challenging due to several factors: (a) Locating the same WBC patch on the peripheral blood film (PBF) across different microscopes and resolutions is inherently difficult [39], (b) Microscope stage scales are not calibrated uniformly across different microscopes, (c) Even within the same microscope, poor alignment leads to resolution inconsistencies, and (d) Viewing a patch at 100x requires immersion oil, whereas 10x and 40x provide better visibility without it. This necessitates repeatedly adding and removing oil when capturing the same patch at multiple resolutions.

### v2

@@ Abstract
Leukemia is the $10^{th}$ most frequently diagnosed cancer and one of the leading causes of cancer-related deaths worldwide. Realistic analysis of leukemia requires white blood cell (WBC) localization, classification, and morphological assessment. Despite deep learning advances in medical imaging, leukemia analysis lacks a large, diverse multi-task dataset, while existing small datasets lack domain diversity, limiting real-world applicability. To overcome dataset challenges, we present a large-scale WBC dataset named ‘Large Leukemia Dataset’ (LLD) and novel methods for detecting WBC with their attributes. Our contribution here is threefold. First, we present a large-scale Leukemia dataset collected through Peripheral Blood Films (PBF) from $48$ patients, through multiple microscopes, multi-cameras, and multi-magnification. To enhance diagnosis explainability and medical expert acceptance, each leukemia cell is annotated at 100x with 7 morphological attributes, ranging from Cell Size to Nuclear Shape. Secondly, we propose a multi-task model that not only detects WBCs but also predicts their attributes, providing an interpretable and clinically meaningful solution. Third, we propose a method for WBC detection with attribute analysis using sparse annotations. This approach reduces the annotation burden on hematologists, requiring them to mark only a small area within the field of view. Our method enables the model to leverage the entire field of view rather than just the annotated regions, enhancing learning efficiency and diagnostic accuracy. From diagnosis explainability to overcoming domain-shift challenges, the presented datasets can be used for many challenging aspects of microscopic image analysis. The datasets, code, and demo are available at: https://im.itu.edu.pk/sparse-leukemiaattri/.
@@
Explainable diagnosis, Blood cancer, leukemia, attribute, sparse annotation, multi-task learning
@@ 1 Introduction
Leukemia is among the leading causes of cancer-related deaths worldwide, with an estimated 487,000 new cases and 305,000 deaths reported in 2022 Bray et al. (2024). This disease disproportionately affects people under the age of 40, with children being particularly impacted Creutzig et al. (2018). The diagnosis of leukemia requires an expert hematologist with extensive experience, capable of analyzing peripheral blood slides employing costly medical equipment. Furthermore, morphological examination remains a cornerstone of leukemia diagnosis, providing accessible, low-cost, and widely available insights into hematologic malignancies Greer and Means (2009). Given the prerequisites for the diagnosis and prognosis of leukemia, many remote regions in underdeveloped countries experienced higher mortality rates Rugwizangoga et al. (2022). However, early detection of leukemia enables timely and accurate treatment, which can significantly improve survival rates. For leukemia diagnosis, hematologists examine the peripheral blood film (PBF) of patients, a process that is often tedious, error-prone, and requires significant expertise. The limited availability of experts and costly diagnostic resources hinders early, precise, and timely detection of leukemia, particularly in resource-constrained areas. Furthermore, leukemia treatment requires repetitive blood sample preparation and detailed analysis, requiring frequent access to experts. AI-driven solutions are vital to overcome the low accessibility of experts. Such a system should be able to work across multiple microscopes and digitization scenarios, and to augment hematologists, it should provide explainable information to hematologists. Unfortunately, existing leukemia datasets  Scotti (2005); Labati et al. (2011); Rezatofighi and Soltanian-Zadeh (2011); Matek et al. (2019); Kouzehkanan et al. (2022); Bodzas et al. (2023) suffer from many limitations, including the absence of localization annotations of white blood cells, the lack of associated morphological attributes, and a limited number of images Gao et al. (2023). Furthermore, there is no established multi-task learning (MTL) He et al. (2024); Xu et al. (2023); Koyuncu et al. (2020) approach that simultaneously detects white blood cells and predicts their morphological attributes.
@@
In addition to these limitations, these supervised algorithms (Sun et al. (2021); Tian et al. (2019); Zhang et al. (2022); Ultralytics (2021); Wollmann and Rohr (2021)) require every cell in each image to be annotated, adding a significant manual burden. In the practical scenario, a hematologist uses only a small portion of the acquired image to perform the required analysis and annotation. Thus majority of the cells remain unannotated. On the other hand, requiring annotation of complete blood films not only increases time and cost but also limits the number of blood films that could be annotated  [^1] Note that multiple image patches are extracted at the specified resolution. Employing existing object detection strategies over sparsely annotated datasets results in subpar performance since these methods are unable to use the unlabeled parts of slides. Complete annotation of a blood film in the monolayer takes around 200 minutes—eight times longer than the 25 minutes needed for sparse annotation.
@@
To overcome dataset challenges, we present a large-scale WBC dataset named ‘large leukemia dataset’ (LLD), that consists of two subsets ‘LeukemiaAttri’ [^2] This subset was published in ‘A Large-scale Multi Domain Leukemia Dataset for the WBC Detection with Morphological Attributes for Explainability’ at MICCAI 2024 Rehman et al. (2024), and this journal paper serves as its extension., capturing cell localization, classification, and morphological attribute prediction tasks; and ‘Sparse-LeukemiaAttri’, for the object-detection and attribute prediction based on sparsely annotated dataset task. An important highlight of the LeukemiaAttri  is that it has been collected through two microscopes from different price ranges, one high-cost (HCM) and one low-cost (LCM), at three magnifications (100x, 40x, 10x) using various sensors: a high-end camera for HCM, a mid-range camera for LCM, and a mobile phone camera for both microscopes. In this collection, using high-resolution HCM (100x), experienced hematologists annotated 10.3k WBC of 14 types, having 55k morphological labels from 2.4k images of several PBS leukemia patients. The data set collection procedure is illustrated in Fig 2 and further details are in section 3.
@@
In conjunction with these datasets, we propose two models named leukemia attribute detector (AttriDet) and sparse Leukemia attribute detector (SLA-Det). For the task of detection of white blood cells (WBCs) along with the prediction of morphological attributes, we integrated a lightweight head, named AttriHead, into the existing object detector. AttriHead employs multiscale features from the feature backbone used in object detection. This not only enhances computational efficiency but also enables the backbone to become more informative through MTL. For the sparsely-supervised object detection, we propose a novel algorithm called SLA-Det. Note that in our case, on average, only 20% of the image is annotated shown in Figure 1. To effectively use the unannotated region information, SLA-Det incorporates region-guided masking and pseudo-label selection based on multiple criteria, ensuring robust model performance. We believe that these extensions align our dataset and framework more with clinical practices and enhance the reliability of automated leukemia detection. Our extensive experimental results and discussion validate the proposed ideas and framework. Our contributions are given below.
@@
    Key results: Using sparse annotations, SLA-Det can process slides at 40x resolutions while maintaining accuracy, unlike AttriDet, which only worked on 100x resolution; thus, it can process slides 4x faster than AttriDet. SLA-Det outperforms state-of-the-art sparse detection methods by 34.7 $mAP@_{50}$, demonstrating superior efficiency and performance.
@@ 2 Related Work
Supervised WBC detection: Deep learning (DL) techniques have been extensively applied to WBC classification Jung et al. (2019); Wu et al. (2023); Wang et al. (2022); Gehlot et al. (2020, 2021). Huang et al. Huang et al. (2020) implemented a two-step approach for their end-to-end solution: first, detecting the WBCs, and then performing classification after cropping the detected cells. To classify single-cell images, Saidani et al. Saidani et al. (2024) optimized a CNN model and employed multiple pre-processing techniques, such as color conversion and image histogram equalization. Similarly, the authors in Thanh et al. (2018) classified two types of WBCs from the ALL-IDB data set Thanh et al. (2018). G Liang et al. Liang et al. (2018) developed a combined model of convolutional neural networks (CNN) and recurrent neural networks (RNN) to classify four types of WBCs using the BCCD dataset Shenggan et al. (2018). To detect WBCs in whole microscopy images, the authors in  Han et al. (2023) gave a lighter deep-learning CNN model and adapted the YOLOX-nano Ge et al. (2021) baseline to identify the types of WBC. Their method enhances the overall model architecture by improving Intersection over Union (IoU) loss. A blood film and bone marrow aspirates-based multiple instance learning for leukocyte identification approaches have been proposed to classify different types of Leukemia  Manescu et al. (2023). For image preprocessing, they mainly applied threshold methods.
@@
Sparsely annotated object detection: To tackle the challenge of object detection in sparsely annotated datasets, authors in Suri et al. (2023) proposed a method that integrates supervised and self-supervised learning into their framework. The approach involves extracting features from an image and its multiple augmented versions, which are then processed through a shared region proposal network. A positive pseudo-mining technique is introduced to effectively filter out labeled, unlabeled, and foreground regions. This results in separate losses for labeled regions, supervised loss and self-supervised loss, ensuring consistent feature extraction across both original and augmented views. This combined method enhances the model’s generalization ability and improves object detection performance. Similarly, Z. Wu et al. Wu et al. (2018) used a soft sampling technique to mitigate the impact of missing annotations. Their approach applies the Gompertz function to adjust gradient weights based on the overlap between positive instances and regions of interest. Niitani et al. Niitani et al. (2019) propose a part-aware learning approach that trains a model twice, using pseudo labels from the first model to guide the second. By leveraging spatial relationships, it selectively ignores classification loss for part categories within subject categories, improving training and performance with sparse annotations. To generate labels for sparsely annotated datasets, the co-mining approach is applied by the authors of Wang et al. (2021). It involves constructing a Siamese detection network with two branches, each generating pseudo-label sets using a co-generation module. It is subsequently used as guided signals to the model, leveraging augmentation techniques to generate diverse pseudo-label sets. Finally, a co-generation module converts predictions from one branch into pseudo-labels for the other, facilitating complementary supervision between the network branches. Note that previous benchmarks, such as Suri et al. (2023); Wang et al. (2021), implemented sparse object detection algorithms on self-created datasets. In contrast, our dataset consists of small patches annotated with detailed morphological features and patch localization. In addition, we introduce a multi-headed WBC detection method tailored for such data.
@@ 3 Dataset
Leukemia comprises four main types: acute lymphocytic leukemia (ALL), chronic lymphocytic leukemia (CLL), acute myeloid leukemia (AML), and chronic myeloid leukemia (CML) Bray et al. (2024). These types are categorized into two major lineages: myeloid (AML and CML) and lymphoid (ALL and CLL). Each leukemia type exhibits distinct abnormalities, including increased myeloblasts in CML, atypical lymphocytes in CLL, elevated myeloblasts in AML, and excessive lymphoblasts in ALL. However, as shown in Table 1, most existing white blood cell (WBC) datasets were collected using a single microscope, lacked bounding boxes and morphological details, and often included either only healthy individuals or patients with a single leukemia type.
@@
Multi. Multi. Cells Bounding Multi. No. of WBC Morphology
Dataset Type Micro. in image Box Res. WBC’s Classes
IDB Scotti (2005) ALL 510 (LB) 2
IDB2 Labati et al. (2011) ALL 260 2
LISC Rezatofighi and Soltanian-Zadeh (2011) Normal 250 5
AML MatekMatek et al. (2019) AML 18,365 15
Raabin Kouzehkanan et al. (2022) Normal 17,965 5
HRLS Bodzas et al. (2023) Multi. 16,027 9
Acevedo Acevedo et al. (2020) Normal 17,092 10
AML Hehr Hehr et al. (2023) AML 81,214 4
WBCAtt Tsutsui et al. (2023) Normal 10,298 5
@@ 3.1 Limitations of existing WBC datasets
The details of publicly available datasets on WBC and leukemia are shown in Table 1, which highlights various limitations such as: 1) The ALL-IDB dataset Scotti (2005) contains only 510 WBCs from patients with Acute Lymphocytic Leukemia (ALL) and includes only cell centroid information. 2) In version 2 of the ALL IDB Labati et al. (2011), each image contains only a single WBC, with the cell’s status labeled accordingly. 3) The LISC Rezatofighi and Soltanian-Zadeh (2011) dataset contains a smaller number of WBCs with five classes, such as basophil, eosinophil, lymphocyte, monocyte, and neutrophil. 4) The entire dataset of AML Matek Matek et al. (2019) and Acevedo Acevedo et al. (2020) contains single-cell images that come under the umbrella of AML and normal individuals. 5) Only Raabin Kouzehkanan et al. (2022), out of all the aforementioned existing datasets, has complete bounding box-level annotation with single-cell images, but the dataset consists entirely of normal individuals, with over 60% of the images containing only a single cell in the FoV. Furthermore, the Raabin-WBC collection contains images from two separate cameras and two different microscopes, without maintaining any correspondence across microscopes. 6) The HRLS Bodzas et al. (2023) dataset is captured using a single microscope and covers only two types of Leukemia and nine types of WBCs. 7) WBCAtt dataset Tsutsui et al. (2023) is the improved version of Acevedo et al. (2020), where annotations of attributes are added for each WBC. The dataset Acevedo et al. (2020) lacks the abnormal WBCs and is captured using a single microscope only.
Figure 2: a) Illustrates the image acquisition procedure utilizing a standard mobile camera (C1), a premium camera for high-cost microscopy (HCM), and a mid-range camera (C2) for low-cost microscopy (LCM) at various resolutions (100x, 40x, and 10x). The images are obtained with both high and low-cost microscopes. b) Displays our microscope set up with four cameras, C1 to C4. C1 and C2 capture the field of view of the microscope lens, and C3 and C4 capture the stage scale. c) This table presents various types of white blood cells (WBC) and their distinctive morphological features.
@@ 3.2 Our dataset collection procedure
To understand leukemia prognosis, we consulted healthcare professionals to finalize key WBC types and their morphology Greer and Means (2009). To improve deep learning across modalities, we propose capturing images of four leukemia PBF types from multiple patients at various magnifications using available resources. Capturing images at varying resolutions enhances deep learning generalization, enabling low-cost, automated leukemia diagnostics.
The dataset comprises 48 Peripheral Blood Films (PBFs) collected from distinct leukemia patients (18 ALL, 22 AML, 2 CLL, 4 CML, and 2 APML), with patient-level splits applied to prevent data leakage and ensure robust evaluation. Images are captured using two microscopes: the high-cost Olympus CX23 and the low-cost XSZ-107BN, along with three cameras: the HD1500T (HCM), ZZCAT 5MP (LCM), and the Honor 9x Lite mobile camera. The high-end camera (ZZCAT 5MP) is 47 times more expensive than the low-end camera (XSZ-107BN), while the HCM is 17 times more costly than the LCM. Although the LCM is more affordable, it sacrifices visual detail and has a smaller field of view. Note that capturing the same WBC across multiple microscopes and resolutions is highly challenging due to several factors: (a) Locating the same WBC patch on the peripheral blood film (PBF) across different microscopes and resolutions is inherently difficult Sultani et al. (2022), (b) Microscope stage scales are not calibrated uniformly across different microscopes. (c) Even within the same microscope, poor alignment leads to resolution inconsistencies, and (d) Viewing a patch at 100x requires immersion oil, whereas 10x and 40x provide better visibility without it. This necessitates repeatedly adding and removing oil when capturing the same patch at multiple resolutions.